# EnsemW2S: Can an Ensemble of LLMs be Leveraged to Obtain a Stronger LLM?

### v1

@@ 3.2 Improving AdaBoost for Complex Generation Tasks
12:     Update $D_{t+1}(i,j)\leftarrow\frac{1}{Z_{t}}D_{t}(i,j)e^{\alpha_{t}\mathbbm{1}\{h_{\theta}^{t}(\bm{x}^{e}_{i},\bm{y}_{i}^{e,j-1})\neq\bm{y}_{i}^{e,j}\}}$ for all $i\in[m],j\in[k_{i}]$ , where $Z_{t}$ is a normalization factor such that $\sum_{i=1}^{m}\sum_{j=1}^{k_{i}}D_{t+1}(i,j)=1$
13:for $o\leftarrow 1$ to $O$ do
14:     for $j\leftarrow 1$ to $k_{o}$ do
15:         Autoregressively generate the $j^{\text{th}}$ token of the “pseudo-answer” $\widehat{\bm{y}}_{o}^{h,j}\sim\Delta^{\text{vocab}}(\sum_{t=1}^{T}\alpha_{t}\cdot\operatorname{softmax}(h^{t}_{\theta}([\bm{x}_{o}^{h},\widehat{\bm{y}}_{o}^{h,1:j-1}])))$, where $\Delta^{\text{vocab}}$ denotes the simplex on the vocabulary
16:Train weak-to-strong student model $f_{\phi}(\cdot)$ on $\{(\bm{x}_{o}^{h},\widehat{\bm{y}}_{o}^{h})\}_{o=1}^{O}$
Weighted Error Calculation. Our weighted error equation $\epsilon_{t}$ also undergoes minor changes. The strict condition for each round of AdaBoost training is that the weighted model error (calculated by comparing each token of each sample) must be less than the pre-training error, i.e., $\epsilon_{t}<\epsilon_{pre}$. The weighted model error $\epsilon_{t}$ is defined as, $\epsilon_{t}=\sum_{i=1}^{m}\sum_{j=1}^{k_{i}}\mathbbm{1}\{h^{t}_{\theta}(\bm{x}^{e}_{i},\bm{y}_{i}^{e,j-1})\neq\bm{y}_{i}^{e,j}\}D_{t}(i,j)<\epsilon_{pre}.$ Here, $\bm{y}_{i}^{e,j-1}$ is the $(j-1)^{\text{th}}$ ground-truth token in the answer part. The model $h^{t}_{\theta}(\bm{x}_{i}^{e},\bm{y}_{i}^{e,j-1})$ predicts the next token and compares it with the ground-truth token $\bm{y}_{i}^{j}$.
Weight Update Equation. Our weight update equation for each token is $D_{t+1}(i,j)\leftarrow\frac{1}{Z_{t}}D_{t}(i,j)e^{\alpha_{t}\mathbbm{1}\{h^{t}_{\theta}(\bm{x}_{i}^{e},\bm{y}_{i}^{e,j-1})\neq\bm{y}_{i}^{e,j}\}}$ where $Z_{t}$ is a normalization factor calculated by taking the norm of the updated weight vector to ensure $\sum_{i=1}^{m}\sum_{j=1}^{k_{i}}D_{t+1}(i,j)=1$.
Combining Experts to Generate Pseudo Answers for Hard Questions: To combine the outputs of different experts trained during the various EnsemW2S-AdaBoost rounds, we scale the probability distribution for each token generated by the model $h^{t}_{\theta}$ in round $t$ by its corresponding weight $\alpha_{t}$. Specifically, we multiply $\alpha_{t}$ by the probability distribution vector of each token. We then aggregate these weighted distributions across all rounds, normalizing the resulting vector to form a new probability distribution for each token.
Using this aggregated distribution, we sample the final predicted token. The process is autoregressive, where the $j^{\text{th}}$ token of the "pseudo-answer" is generated as
@@
$\widehat{\bm{y}}_{o}^{h,j}\sim\Delta^{\text{vocab}}\left(\sum_{t=1}^{T}\alpha_{t}\cdot\operatorname{softmax}\left(h^{t}_{\theta}\left([\bm{x}_{o}^{h},\widehat{\bm{y}}_{o}^{h,1:j-1}]\right)\right)\right)$ (1)
@@
By combining the outputs of multiple experts, each trained in different EnsemW2S-AdaBoost rounds, the ensemble approach leverages diverse perspectives from the weak models. Each expert contributes its learned strengths, and through weighted aggregation, we diminish the influence of models that are less confident or less effective on certain tokens. This helps reduce variance in the generation process, ensuring that errors from individual weak models are mitigated. The result is a more robust pseudo-labeling system that is better aligned with the true distribution of the hard data, often yielding a performance improvement over any single weak model.
Unlike classification, where scores are combined over a fixed set of classes, generation tasks involve predicting sequences of tokens, where each prediction affects future ones. This makes combining generation probabilities more complex, as errors in early token predictions can propagate throughout the sequence. Additionally, we are aggregating probability distributions over large vocabularies, which introduces computational overhead and potential numerical instability.
Our method addresses these challenges by using a weighted combination of expert models’ token probabilities, ensuring that weaker predictions from individual rounds are minimized. By normalizing the aggregated distribution for each token, we maintain valid probability distributions across the vocabulary, effectively reducing the risk of cascading errors during autoregressive generation. This ensemble approach results in a more stable and accurate generation process, mitigating the issues inherent in sequence modeling.
@@
Pseudo answer generation on multiple-choice datasets: On multiple-choice datasets, instead of using generated tokens $\widehat{\bm{y}}^{h}$ as pseudo answers, we can select one of the choices in the MCQ dataset using negative log-likelihood (NLL). Specifically, we calculate the NLL between the choices and $\widehat{\bm{y}}^{h}$ and select the choice with the lowest NLL. For datasets without multiple choices, we can directly use $\widehat{\bm{y}}^{h}$.
Train W2S Model: The strong student model, $f_{\phi}(\cdot)$, is trained using pseudo answers generated for the hard data $\{(\bm{x}_{o}^{h},\widehat{\bm{y}}_{o}^{h})\}_{o=1}^{O}$. While it might be beneficial to include the labeled easy data in the training process, we adhere to the pipeline established by Burns et al. (2023) by focusing exclusively on the hard examples to maintain consistency.
@@
Ablation Studies. We experimented with combining the logits directly instead of probabilities but did not observe any improvement (refer to Appendix Figure 11). We conducted ablation studies where, instead of treating each token as independent, we used a sliding window of length $L$ while calculating weights and aggregating errors (see Appendix Figure 12 and 13). Different window lengths did not cause significant changes in values, so we ultimately chose a window of $L=1$. We also explored treating each sample as independent instead of each token as independent in the sample-answer part, finding better results with the latter. This is reasonable since the error calculated using independent-sample weights is less accurate.
@@
Evaluation Metric. We used two metrics to evaluate this Q/A dataset. One is (1) Token-wise comparison, where we compare each predicted token and average the total error, and (2) Option-wise comparison, where we compare the negative log-likelihood (NLL) of the correct answer completion with the NLLs of the incorrect answer completions. Accuracy represents the number of entries where the correct answer completion has the lowest NLL among all choices.
@@ 4 Related Work
Weak-to-Strong (Burns et al., 2023) was the first to introduce the problem of weak-to-strong generalization for the super-alignment problem, where the ultimate aim is to elicit the full capabilities of the strong model using supervision only from weak models. (Charikar et al., 2024) provides a theoretical framework for the same with insights on how much w2s improvement can occur, though their work is limited to a few layer neural networks. Similarly, (Lang et al., 2024) provides bounds on expansion properties using finite data distributions for when w2s generalization will happen, but only for simple binary classification tasks. (Zhang et al., 2024) proves that transcendence (exceeding the capability of the model that generates the training data) is possible for low-temperature sampling. Although this setting is not exactly w2s, it sheds light on this direction.
Several works have attempted to solve w2s generalization in LLMs. (Sang et al., 2024) tries to improve this supervision using ensemble learning and scalable oversight for binary classification NLP tasks but cannot observe significant improvement. (Ji et al., 2024) introduces a model that enhances the alignment of LLMs with human intentions by correcting the residual differences between aligned and unaligned answers by training on a query-answer correction dataset. This method boosts w2s generalization using supervisory signal from smaller models to improve the performance of complex systems. In (Sun et al., 2024), the authors propose a scalable approach for e2h generalization which involves training reward models on easier tasks and using them to evaluate performance on harder tasks. (Liu & Alahi, 2024) introduces a method similar to the classical hierarchical mixture of experts, where multiple specialized weak supervisors are used for weak-to-strong generalization instead of a single generalist model. (Bansal et al., 2024) compares large LLM training from data generated using weak (cheap) vs strong (expensive) model in a compute matching way and finds larger data from weaker model to provide better w2s.
Ensemble Learning Binary Classification Boosting (Freund & Schapire, 1997) and multi-classfication boosting (Hastie et al., 2009) are common ensemble learning algorithms. In (Verga et al., 2024), they use a voting mechanism to combine multiple small LLMs instead of a single large LLM to evaluate another LLM and show it performs better than large LLMs. An extended related work section is present in Appendix A.
@@ 5 Experimental Setup
We test two different strategies for each task. One aligns with Burns et al. (2023), where we split the training data randomly into train-weak and train-strong. Train-weak is used to train the weak model. Train-strong is used to train the strong and transfer models using pseudo labels generated using the weak model. The second strategy involves splitting the training data into easy and hard splits, where the easy data is now train-weak, and the hard data is now train-strong with the same training pipeline. This is also a more realistic setup for weak-to-strong generalization, as discussed in Section 1. For both strategies, we aim to recover the performance gap (PGR) and elicit the full capability of the strong model using an ensemble of weak models. The baseline in all experiments uses a single model for w2s generalization, following the principle of Burns et al. (2023).
We run AdaBoost/EnsemW2s-AdaBoost 10 times for the binary classification tasks and 5 times for the generation tasks. We pick the best w2s performing round for our plots. However, we observe that all rounds ($n>=2$) are better than the baseline ($n=1$). Additionally, we chose single model performance ($n=1$) for weak model performance.
5.1 Binary Classification Task
Figure 2: Binary Classification Task: Top figure shows a bar plot comparing w2s generalization of our method (grey) with a baseline (blue) from Burns et al. (2023) using accuracy values(%) for different combinations of weak and strong model pairs for random data split (top bar-plot) and easy-hard split(bottom bar-plot). Bottom figure shows a line plot comparing the accuracy and performance gap recovered values (PGR). The left two figures are for random data split, while the right two figures are for the easy-hard split to show e2h generalization.
Scaling Law: In Figure 2 (line plot), we see less PGR recovery for the Qwen-1.8B model even though it is similar in size to GPT-2-xl. Similarly, in the bar plot, we see a drastic difference between the oracle performance of GPT2xl and Qwen-1.8B. This is because the Qwen models series are more capable even after being the same size. Thus, model size is not a good metric, but model capability is a better metric for differentiating between weak and strong models.
Better metric: Figure 2 shows the accuracy and PGR plots for both random and easy-hard split. We observe that PGR is not very informative, as it can produce extremely large or even negative values. In the w2s experiments, large values occur because the ensemble of weak models becomes strong enough to match or exceed a strong model, improving w2s generalization. Negative values, seen in baseline experiments, indicate the transfer model performed worse than the weak model, often when the strong model fails to learn and its inductive bias becomes random with pseudo-label training. Similar patterns are seen in Figure 5 and 4. (Refer to Appendix Table 1 and 2 for more details.)
5.2 Generation Task for Multiple Choice Dataset
5.2.1 Comparing Weak model’s performance
Figure 3: Performance comparison of a single weak model (dark color) with the combined weak models (Lighter hue shows improvement).
In Figure 3, we compare the performance of a single weak model (dark color) with combined weak models after 5 rounds of EnsemW2S-AdaBoost. Smaller models show greater improvement, which is expected since boosting works best when weak models are diverse. Using EnsemW2S-AdaBoost, smaller models can diversify through the data sampling step; however, larger models tend to learn all possible information and cannot learn something different with each round. Also, we use token error in Figure 3 since it is a more precise metric to measure improvement in weak models.
5.2.2 Comparing Strong model’s performance
Here, we use the multiple-choice classification accuracies to calculate the accuracy of all our plots. We show the accuracy values of token-wise metrics in the Appendix tables.
W2S Results with Random Training Data Splits. From Figure 4 and 5, we see that w2s training using an ensemble of experts almost consistently outperforms the baseline (single expert). Thus, ensemble learning is beneficial. We can see the trend of accuracy and performance gap recovered for the different model pairs in Figure 4 and 5 for Quartz and ARC datasets, respectively. For Quartz data, we see that our PGR percentage (Figure 4) improves as the model scales up except when the weak model is the smallest sized model (pythia-70m). This could be because the increasing capability difference between the small and large models makes it difficult for the strong model to learn anything from the weak. This trend is the same in the baseline as well as our EnsemW2S. But an important thing to note is that for some cases for both ARC and Quartz data, our method generates a large PGR percentage of >=100%, showing the ability of our w2s method to recover the performance gap.
W2S Results with Easy-Hard Training Data Splits. From Figure 4 and 5, we see that w2s training using an ensemble of experts almost consistently outperforms the baseline (single expert). Thus showing that ensemble learning is beneficial. Our method shows more improvement over baseline for easy-hard data split as compared to random split. This is because of two reasons. Firstly, the power of combining weak models using our modified AdaBoost is more useful when all of them are weak but slightly different from each other. Secondly, by easy and hard splitting, the margin between weak and strong increases more, giving more room for improvement.
We also observe that PGR for e2h generalization is significantly lower, highlighting the complexity of the e2h generalization problem. We hope this work could motivate researchers to build more sophisticated methods for this more complex e2h generalization problem. Another simple observation is as the models become more capable, both the performances (baseline and ours) increase.

### v2

@@ 3.3 EnsemW2S: LLM Token-Level Ensemble Method
10:     Update sample weights: $D_{t+1}(i,j)\leftarrow\frac{D_{t}(i,j)}{Z_{t}}e^{\alpha_{t}\mathbbm{1}\{h_{\theta}^{t}(x_{i}^{e},y_{i}^{e,j-1})\neq y_{i}^{e,j}\}}\forall i\in[m],j\in[k_{i}]$, where $Z_{t}=\sum_{i=1}^{m}\sum_{j=i}^{k_{i}}D_{t}(i,j)\cdot e^{\alpha_{t}\mathbbm{1}(h_{\theta}^{t}(\bm{x}^{e}_{i},\bm{y}_{i}^{e,j-1})\neq\bm{y}_{i}^{e,j})}$
11:  for $o\leftarrow 1$ to $O$ do
12:     for $j\leftarrow 1$ to $k_{o}$ do
13:        Autoregressively generate the $j^{\text{th}}$ token of the “pseudo-answer”
@@
$\hat{y}_{o}^{h,j}\sim\Delta_{\text{vocab}}\bigl{(}\sum_{t=1}^{T}\alpha_{t}\cdot\text{softmax}(h_{\theta}^{t}(x_{o}^{h},\hat{y}_{o}^{h,1:j-1}))\bigr{)},$
where $\Delta^{\text{vocab}}$ denotes the simplex on the vocabulary
14:  Train weak-to-strong model $f_{\phi}(\cdot)$ on $\{(x_{o}^{h},\hat{y}_{o}^{h})\}_{o=1}^{O}$
Incorporating Prior Term. Vanilla Adaboost for binary classification uses $\alpha_{t}=\log({\frac{1-\epsilon_{t}}{\epsilon_{t}}})$ to calculate votes for each classifier. To keep $\alpha_{t}$ positive they introduce a weak learning condition, $\epsilon_{t}<0.5$, which means that weighted error for each classifier should be better than random. However, using the same weak-learning condition and voting mechanism is not feasible for multi-classification tasks. Thus, Hastie et al. (2009) uses an additional prior $\log(c-1)$ term, where $c$ is the number of classes, in the calculation of the AdaBoost parameter $\alpha_{t}$. This term serves two purposes: (1) It enables the generation of weak models with accuracy above $\frac{1}{c}$, where $\frac{1}{c}$ is random selection accuracy. This is crucial for smaller models and challenging tasks that cannot achieve 50% accuracy. (2) It ensures that $\alpha$ remains positive. Now, for generative LLMs with large vocabularies, $\log(c-1)$ renders $\alpha_{t}$ nearly identical, making this approach infeasible. Therefore, we introduce a different method for calculating votes, $\alpha_{t}\leftarrow\log({\frac{1-\epsilon_{t}}{\epsilon_{t}}})+\log(\frac{1}{1-\epsilon_{pre}}-1)$, where $\epsilon_{pre}$ is the pre-trained model error of the chosen LLM. We provide a theoretical proof (Theorem 1 in Section 3.3) regarding how the training error bound of the ensemble weak learners reduces exponentially with the addition of new weak learners while using new $\alpha_{t}$ method. This term is sensible because $\epsilon_{pre}$ represents the pretrained model error and makes our new weak learning condition to be $\epsilon_{t}<\epsilon_{pre}$, effectively replacing the binary and multi-class AdaBoost’s conditions.
Weighted Error Calculation. Following the new weak learner’s condition, the strict condition for each weak expert training is now that the weighted model error (calculated by comparing each token of each sample) must be less than the pre-training error, i.e., $\epsilon_{t}<\epsilon_{pre}$. The weighted model error $\epsilon_{t}$ is defined as, $\epsilon_{t}=\sum_{i=1}^{m}\sum_{j=1}^{k_{i}}\mathbbm{1}\{h^{t}_{\theta}(\bm{x}^{e}_{i},\bm{y}_{i}^{e,j-1})\neq\bm{y}_{i}^{e,j}\}D_{t}(i,j)<\epsilon_{pre}.$ Here, $\bm{y}_{i}^{e,j-1}$ is the $(j-1)^{\text{th}}$ ground-truth token in the answer part. The model $h^{t}_{\theta}(\bm{x}_{i}^{e},\bm{y}_{i}^{e,j-1})$ predicts the next token and compares it with the ground-truth token $\bm{y}_{i}^{j}$ (during training when ground-truth is still available).
Weight Update Equation. Our sample-weight update equation for each token is $D_{t+1}(i,j)\leftarrow\frac{1}{Z_{t}}D_{t}(i,j)e^{\alpha_{t}\mathbbm{1}\{h^{t}_{\theta}(\bm{x}_{i}^{e},\bm{y}_{i}^{e,j-1})\neq\bm{y}_{i}^{e,j}\}}$ where $Z_{t}$ is a normalization factor ensuring that the updated weights satisfy $\sum_{i=1}^{m}\sum_{j=1}^{k_{i}}D_{t+1}(i,j)=1$. The main idea is to adjust the sample weights to emphasize misclassified examples, thereby guiding the sampling process for next weak learner.
Combining experts to Generate Pseudo Answers for Hard Questions. To combine the outputs of different experts trained during the various EnsemW2S rounds, we scale the probability distribution for each token generated by the model $h^{t}_{\theta}$ in round $t$ by its corresponding weight $\alpha_{t}$. Specifically, we multiply $\alpha_{t}$ by the probability distribution vector of each token. We then aggregate these weighted distributions across all rounds, normalizing the resulting vector to form a new probability distribution for each token. Using this aggregated distribution, we sample the final predicted token. The process is autoregressive, where the $j^{\text{th}}$ token of the “pseudo-answer” is generated as
$\widehat{\bm{y}}_{o}^{h,j}\sim\Delta^{\text{vocab}}\left(\sum_{t=1}^{T}\alpha_{t}\cdot\operatorname{softmax}\left(h^{t}_{\theta}\left([\bm{x}_{o}^{h},\widehat{\bm{y}}_{o}^{h,1:j-1}]\right)\right)\right)$
@@
By combining the outputs of multiple experts, each trained in different EnsemW2S rounds, the ensemble approach leverages diverse perspectives from the weak models. Each expert contributes their learned strengths, and through weighted aggregation, we diminish the influence of models that are less confident or less effective on certain tokens. This helps reduce variance in the generation process, ensuring that errors from individual weak models are mitigated. The result is a more robust pseudo-labeling system that is better aligned with the true distribution of the hard data, often yielding a performance improvement over any single weak model.
@@
Pseudo Answer Generation on Multiple-Choice Datasets. On multiple-choice Q/A datasets, instead of using generated tokens $\widehat{\bm{y}}^{h}$ as pseudo-answers, we can select one of the choices in the MCQ dataset using negative log-likelihood (NLL). Basically, we calculate the NLL between the choices and $\widehat{\bm{y}}^{h}$ and select the choice with the lowest NLL. For datasets without multiple choices, we can directly use $\widehat{\bm{y}}^{h}$.
How does EnsemW2S improve generalization? First, Theorem 1 (3.3) shows that the training error decreases as more weak learners are added. Next, the “Margin theory for generalization” demonstrates that the margin distribution over the training data improves with each additional weak learner. Consequently, adding new models in the EnsemW2S ensemble reduces training error without overfitting, thereby improving generalization.
Theorem 1 (Exponential decay of training error with additional weak learners) Under the condition $\epsilon_{t}<\epsilon_{\text{prev}}$, our token-level ensembling method, EnsemW2S, reduces the training error exponentially:
$\text{Error}_{\text{train}}=\frac{1}{n}\sum_{i=1}^{m}\sum_{j=1}^{k_{i}}\sum_{t=1}^{T}\alpha_{t}\mathbbm{1}\left(h_{\theta}^{t}(\bm{x}^{e}_{i},\bm{y}_{i}^{e,j-1})\neq\bm{y}_{i}^{e,j}\right)<\prod_{t=1}^{T}Z_{t},$
where $Z_{t}<1$, $T$ is the number of EnsemW2S rounds, $n=\sum_{i=1}^{m}k_{i}$, and $k_{i}$ is the token length of the $i^{\text{th}}$ easy example $S^{e}=\{(\bm{x}^{e}_{i},\bm{y}^{e}_{i})\}_{i=1}^{m}$. We assume tokens are i.i.d. for this proof. See Appendix F for full details.
Margin Theory for Generalization. Drawing on results from AdaBoost (binary classification) and SAMME (multiclass boosting), we observe that the margin over training data increases with the number of weak learners, making the ensemble’s predictions more confident. Prior work shows increased margins improve generalization. Since we assume i.i.d. tokens, these conclusions can be directly apply. See Appendix G for details.
Train W2S Model. The strong student model, $f_{\phi}(\cdot)$, is trained using pseudo answers generated for the hard data $\{(\bm{x}_{o}^{h},\widehat{\bm{y}}_{o}^{h})\}_{o=1}^{O}$. Unlike Burns et al. (2023), we also include easy human-level data in the training, as it is a sensible choice. For completeness and proper comparison, we additionally run experiments that exclude the easy data, following Burns et al. (2023).
@@
Ablation Studies. Logit combination instead of probabilities showed no improvement (see Appendix Figure 9). We conducted ablation studies where, instead of treating each token as independent, we used a sliding window of length $L$ while calculating weights and aggregating errors (see Appendix Figure 10 and 11). Different window lengths did not cause significant changes in values, so we ultimately chose a window of $L=1$. We also explored treating each sample as independent instead of each token as independent in the sample-answer part, finding better results with the latter. This is reasonable since the error calculated using independent-sample weights is less accurate.
@@
Evaluation Metric. We used two metrics to evaluate this Q/A dataset. One is (1) Token-wise comparison, where we compare each predicted token and average the total error, and (2) Option-wise comparison, where we compare the negative log-likelihood (NLL) of the correct answer completion with the NLLs of the incorrect answer completions. Accuracy represents the number of entries where the correct answer completion has the lowest NLL among all choices. We use option accuracy to reported results in the main paper.
Computational cost of EnsemW2S. The primary increase in computational cost compared to the baseline—where a single weak expert supervises the student—comes from multi-LLM decoding. However, its important to consider that unlike other multi-LLM works Anonymous (2025); Du et al. (2023), we do not use an additional reward or judge model for combining multi-LLM outputs, significantly reducing hardware demands. Moreover, we rely on the same labeled data rather than collecting new data to train separate models. A detailed computational analysis of each EnsemW2S component is provided in Appendix I.
@@ 4 Related Work
Weak-to-Strong. (Burns et al., 2023) first introduced w2s generalization for super-alignment, aiming to elicit strong model capabilities using only single weak model supervision. (Charikar et al., 2024) provides a theoretical framework for the same with insights on how much w2s improvement can occur, though their work is limited to a few layers of neural networks. (Lang et al., 2024) establishes expansion bounds for w2s generalization under finite data distributions, but focuses solely on binary classification. (Zhang et al., 2024) shows that transcendence—surpassing the capability of the training data source—is possible via low-temperature sampling, offering insights relevant to w2s.
Several works have attempted to solve w2s generalization in LLMs. (Sang et al., 2024) explores ensemble learning for and scalable oversight for binary classification NLP tasks, but sees limited gains. (Ji et al., 2024) introduces a model that enhances the alignment of LLMs with human intentions by correcting the residual differences between aligned and unaligned answers by training on a query-answer correction dataset. It enhances w2s generalization by leveraging smaller models for supervisory signals. (Sun et al., 2024) proposes a scalable e2h generalization method, training reward models on simpler tasks and using them to evaluate harder ones. (Liu & Alahi, 2024) adapts the hierarchical mixture of experts, using multiple specialized weak supervisors instead of a single generalist for w2s. (Bansal et al., 2024) compares LLM training on weak (cheap) vs. strong (expensive) model-generated data and finds that larger weak-model datasets yield better w2s.
Ensemble Learning. Binary Classification Boosting (Freund & Schapire, 1997) and multi-classification boosting (Hastie et al., 2009) are common ensemble learning algorithms. In (Verga et al., 2024), they use a voting mechanism to combine multiple small LLMs instead of a single large LLM to evaluate another LLM and show it performs better than large LLMs. An extended related work section is present in Appendix C.
@@ 5 Experimental Setup
ID and OOD Data Setup.
To assess our method’s improvement in generalization on both in-distribution (ID) and out-of-distribution (OOD) datasets, we use factual question-answering datasets such as ARC (Clark et al., 2018) and Quartz (Tafjord et al., 2019), as well as more complex mathematical datasets like math-mc (Tan et al., 2024) and MMLU (categories: high-school mathematics, elementary mathematics, college mathematics) (Hendrycks et al., 2021). All these datasets feature multiple-choice question-answering tasks, facilitating pseudo-label generation based on options.
For studying ID performance, we employ ARC, Quartz, and math-mc datasets, randomly splitting each dataset into two halves. One half is used for training weak experts. We treat the second half as unlabeled, generating pseudo-labels from the trained weak experts. Subsequently, we train the strong student model using both this pseudo-labeled data and the initially labeled half. For assessing OOD performance, we adopt two distinct strategies. First, we partition datasets into easy and hard subsets, training weak experts on labeled easy data, and the strong student on unlabeled hard data combined with labeled easy data (similar to the ID setup). We evaluate performance on a small, labeled subset sharing the hard data distribution. Difficulty levels in math-mc are directly available through question-difficulty ratings, whereas for ARC and Quartz, we generate difficulty ratings via a cross-validation approach described in Appendix H. Difficulty rating plots for ARC and Quartz are provided in Figures 7 and 8, respectively. The second strategy involves training weak and strong models on one dataset (math-mc) and evaluating performance on a different dataset (MMLU’s mathematics categories) to study how cross-data OOD generalization changes.
Expert and Student Models.
We employ Qwen2.5-1.5B and Qwen2.5-3B models as weak experts, and Qwen2.5-7B as the strong student to study both ID and OOD generalization. Additionally, we explore scaling laws through multiple combinations of weak and strong model pairs using the Quartz dataset. For this scaling analysis, we utilize the Pythia model series, ranging from 70M to 2.8B parameters since pythia even though relatively old model series provided us with a wide range of models sizes. To determine the optimal number of AdaBoost rounds—i.e., how many models to combine—we use a validation set matching the training distribution. We limit our experiments to a maximum of 5 AdaBoost rounds/models.
Baseline.
We compare EnsemW2S to a single weak expert and a strong student trained on outputs of single weak expert. Additional baselines include training the strong student on weak experts’ labeled data and strong student’s pre-training performance.
6 Results
6.1 Enhancing ID and OOD Generalization of Weak Experts using EnsemW2S
Figure 3: Performance comparison of weak models (single vs. ensemble). Blue bars: Qwen2.5-3B; orange: Qwen2.5-1.5B. Dark shades show single-model performance; light shades show ensemble gains. The bars are grouped into: (1) ID evaluation (first three subplots), where train and test distributions are similar; (2) OOD via easy-hard splits (middle three), where models are trained on the easy and tested on the hard; and (3) cross-dataset OOD generalization (last three), where models trained on the easy of Math-MC are evaluated on three difficulty levels of MMLU.
As depicted in Figure 3, we analyze the in-distribution (ID) and out-of-distribution (OOD) generalization performance of our EnsemW2S-based combination of weak experts for two models, Qwen2.5-1.5B and Qwen2.5-3B. We compare the performance of a single weak model (darker color) against an ensemble of weak models trained on the same labeled dataset, with lighter hues indicating performance improvements. For ID generalization, we observe absolute improvements between of 0.4-3.02%, 2.3-4%, and 2.24-2.26% on the ARC, Quartz, and Math-MC datasets, respectively. For OOD generalization based on an easy-to-hard data split, the ensemble achieves improvements between 1.02-2.84%, 1.03-1.66%, and 3.7-4.34% on the same datasets. We also evaluate cross-dataset OOD performance by training on the easy split of Math-MC and testing across three difficulty levels in MMLU’s mathematics categories. In this setting, we observe gains between of 2.07-2.7%, 3-6%, and 3-4% for elementary, high school, and college-level math questions, respectively. These results suggest that generalization performance is more sensitive to question difficulty than to dataset variation.
As expected, performance of ID data is higher than to OOD data. However, the consistent gains achieved by our ensemble method—despite using the same labeled data as the single-model baseline—highlight its robustness and effectiveness. A few simple baselines are presented in Appendix J.1.
6.2 Stronger Expert further Improve Student and its W2S generalization performance.
In-Distribution Performance
Qwen2.5-3B Qwen2.5-7B
Dataset Single Expert EnsemW2S Expert Student (Baseline) Student (Ours) Oracle Student on Expert’s data Pre-trained Student
ARC 54.02 57.04 56.74 58.70 59.39 56.04 45.48
Quartz 81.25 85.20 84.95 88.10 88.78 83.47 60.00
Math-mc 61.74 64.00 62.90 65.00 65.87 62.90 24.73
Out-of-Distribution Performance
Qwen2.5-3B Qwen2.5-7B
Single Expert EnsemW2S Expert Student (Baseline) Student (Ours) Oracle Student on Expert’s data Pre-trained Student
ARC 43.00 45.76 45.39 46.67 51.19 45.42 34.00
Quartz 80.99 82.02 83.29 85.71 87.5 82.48 56.12
Math-mc 33.00 37.00 55.86 59.74 60.91 49.00 26.72
MMLU-Elementary 71.43 73.28 73.91 76.19 75.66 73.57 44.71
MMLU-High-Skl 50.00 53.00 52.59 53.70 54.00 52.18 23.70
MMLU-College 33.00 37.00 46.00 47.00 47.50 44.00 33.00
Table 1: The first two columns show performance of a single weak expert and the EnsemW2S based ensemble. Its followed by strong students trained on the single and combined experts. The table also includes an oracle upper bound and two baselines: one trained on weak labels and one showing pretrained performance.
As shown in Table 1, we study how an ensemble of weak models improves the generalization performance of the strong model. We focus on Qwen2.5-7B for this analysis and evaluate both ID and OOD generalization, following the same setup as before. For ID generalization, we observe absolute performance improvement of 2%, 3.2%, and 2.1% on the ARC, Quartz, and Math-MC datasets, respectively. For OOD generalization based on an easy-to-hard split, we observe gains of 1.28%, 2.42%, and 1.17% on the same datasets. We also evaluate cross-dataset OOD performance by training the ensemble on the easy subset of Math-MC and testing on varying difficulty levels within MMLU’s mathematics categories. Here, we observe improvements of 2.28%, 1.11%, and 1% for high-school, elementary, and college-level math questions, respectively.
A few notable conclusions: the EnsemW2S-based ensemble of weak experts achieves performance comparable to the W2S-trained strong student, demonstrating the effectiveness of ensembling. It even outperforms a strong model trained on weak experts’ labeled data, despite the total combined model size being comparable in most cases. Moreover, the strong model trained with weak supervision from the ensemble surpasses the baseline which use a single expert model.
6.3 Improvement in PGR and W2S generalization for all range of model sizes.